When one mole of non-volatile solute is dissolved in three moles of solvent the vapour pressure of the solution relative to the vapour pressure of the pure solvent is?

Grade:12

1 Answers

Prashik Gaikwad
6 years ago
Since Relative lowering vapour pressure is equal to mole fraction of solute i.e., P°/P°1=X2=n2/n1+n2=1/1+3=1/4=0.25 mm HgTherefore Relative lowering vapour pressure of pure solvent in solution is 0.25 mm Hg